OpenStudy (tjhawkes):

A company has two packaging machines in a unit, each with a different daily capacity. The capacity of machine 1 is defined by the function f(m) = (m + 4)2 + 100, and the capacity of machine 2 is defined by the function g(m) = (m + 12)2 − 50, where m is the number of minutes the packaging machine operates. Create the function C(m) that represents the combined capacity of the two machines.

OpenStudy (kevin):

c(m) = f(m) + g(m) c(m) = (m + 4)^2 +100 + (m + 12)^2 − 50 Then simplify